The radius of the earth is about 6370km. An object that has a mass of 20kg is taken to a height of 160km above the earth surface. (a) How much does the object weigh 


1
Expert's answer
2021-04-30T11:23:25-0400
"W=mg\frac{R^2}{(R+h)^2}\\W=(20)(9.8)\frac{6370^2}{(6370+160)^2}=186.5\ N"